OpenStudy (anonymous):

The ellipse has a foci located at (3,1) and (7,1) and a major axis length of 10. What is the length of the minor axis?

OpenStudy (amoodarya):

a=5,c=2 b^2=a^2-c^2=21 so b=root21 look at this image

OpenStudy (anonymous):

@amoodarya thank you, is that the graph of the ellipse? what would the minor axis be in that case?

OpenStudy (amoodarya):